Determine the homology groups of the Klein bottle.
Answer #1

Homology of the klein bottle b v 3,4 = atb-4 da L = ita-c С. id &qu= atb-c gL::20 v Б CL XQ Z o 0 } { 9 ZZ ĐC a b b c a, badoif Pt29. =0 ? or po P=1=0 45037 Koros o generated 2d: and atta Goº(k) In de is КІ до — 7 - 2. Imdi=0 atbic /0 Hiºl) keyal ZOZplease upvote